OpenStudy (anonymous):

Why it is essential that a racetrack turn to be banked so that it stops downhill toward the center of the turn?

OpenStudy (ujjwal):

When the road is banked (say, at an angle \(\theta\)), the normal force acts an angle inclined to vertical. If the normal force be R, the vertical component Rsin\(\theta\) balances the weight of the vehicle and the horizontal component Rcos\(\theta\) provides the necessary centripetal force at the turn..